# HiLo: A Learning Framework for Generalized Category Discovery Robust to Domain Shifts

### v1

@@ 3.2.1 Learning domain-semantic disentangled features for GCD
As covariate shift observed by new domains $\Omega$ in $\mathcal{D}^{u}$ degrades performance, we aim to learn two distinct feature sets encoding domain and semantic aspects by minimizing their mutual information. For each image $\boldsymbol{x}$, we thus consider that its feature can be partitioned into two parts, depicting domain-specific (e.g., real, sketch) and semantic information (e.g., cat, dog), respectively. However, it is intractable to estimate the mutual information between random variables of semantic and domain in finite high-dimensional space without parametric assumptions [63, 64]. Instead of calculating the exact value, assumptions based on convex conjugate [65] and GAN [66] are utilized for estimation. [67, 68] further demonstrate that this estimation can be achieved without such assumptions. We thus adopt the approach from [68] based on Jensen-Shannon divergence to estimate the mutual information. For each image, instead of considering a single feature vector as $\boldsymbol{z}=\mathcal{H}(\mathcal{F}(\varphi(\boldsymbol{x})))$, here we consider two feature vectors, $\boldsymbol{z}_{d}$ and $\boldsymbol{z}_{s}$, for domain and semantic information respectively. Inspired by the fact that deeper layers of the model give higher-level features and the shallower layers of the model give lower-level features [69, 25], we use the feature from the very first layer of the ViT as $\boldsymbol{z}_{d}$ and that from the very last layer as $\boldsymbol{z}_{s}$. Specifically, we obtain $[\boldsymbol{z}_{d},\boldsymbol{z}_{s}]=\tilde{\mathcal{H}}(\mathcal{F}(\varphi(\boldsymbol{x})))$, where $\tilde{\mathcal{H}}$ consists of two projection heads, one on the first layer feature of $\mathcal{F}$ and the other on the last layer feature of $\mathcal{F}$ (see Figure 2). Therefore, the mutual information between domain and semantic features can be approximated by a Jensen-Shannon estimator:
$\mathcal{L}_{m}={I}_{\Phi}(\boldsymbol{z}_{d},\boldsymbol{z}_{s})=\mathbb{E}_{p(\boldsymbol{z}_{d},\boldsymbol{z}_{s})}\left[-\log\left(1+e^{-\Phi(\boldsymbol{z}_{d},\boldsymbol{z}_{s})}\right)\right]-\mathbb{E}_{p(\boldsymbol{z}_{d})p(\boldsymbol{z}_{s})}\left[\log\left(1+e^{\Phi(\boldsymbol{z}_{d},\boldsymbol{z}_{s})}\right)\right],$ (4)
where $\Phi$ is an MLP and an output dimension of 1. $\Phi$ takes the concatenation of $\boldsymbol{z}_{s}$ and $\boldsymbol{z}_{d}$ as input and predict a single scalar value. We aim to minimize the expected log-ratio of the joint distribution concerning the product of marginals. Note that here $\boldsymbol{z}_{s}$ and $\boldsymbol{z}_{d}$ may come from two different images. In practice, we tile the domain and semantic features of all the images in the mini-batch, and concatenate them, before applying $\Phi$ on all the concatenated features. We then extract the diagonal entries (which are from the marginals) as the first term and the other entries (which are from the joint distribution) as the second term in Equation 4.
@@ 3.2.2 PatchMix contrastive learning
Mixup [70] is a powerful data augmentation technique that involves blending pairs of samples and their corresponding labels to create new synthetic training examples. It has been shown to be very effective in semi-supervised learning [71], long-tailed recognition [72], etc. In the presence of domain shifts, Mixup has also been shown to be effective in unsupervised domain adaptation [73] and domain generalization [25]. Recently, PMTrans [17] introduced PatchMix, which is a variant of Mixup augmentation by mixing up the embeddings of images in the Transformer-based architecture for domain adaptation. Particularly, for an input image $\boldsymbol{x}$ with label $y$, PatchMix augments its $j$-th embedding patch by
$\bar{\varphi}(\boldsymbol{x})_{j}={\beta}_{j}\odot\varphi(\boldsymbol{x})_{j}+(1-{\beta}_{j})\odot\varphi(\boldsymbol{x}^{\prime})_{j},$ (5)
@@
where $\boldsymbol{x}^{\prime}$ is an unlabelled image with or without domain shift, ${\beta}_{j}\in[0,1]$ is the random mixing proportion for the $j$-th patch, sampled from Beta distribution, and $\odot$ denotes the multiplication operation. A one-hot vector derived from $y$ is then smoothed based on $\beta_{j}$ to supervise the cross-entropy loss to train the classification model. However, this works under the assumption that the out-of-domain samples share the same class space with the in-domain samples, restricting its application to the more practical scenarios where the out-of-domain samples may come from new classes as we consider in the problem of GCD with domain shift. Hence, we devise a PatchMix-based contrastive learning method to address the challenge of GCD in the presence of domain shift. Our approach properly leverages all available samples, including both labelled and unlabelled data, from both in-domain and out-of-domain sources, encompassing both old and new classes. By incorporating these diverse samples, our technique aims to improve the model’s ability to handle domain shifts and effectively generalize across different classes.
@@
When incorporating the PatchMix into our problem setting, the unlabelled sample $\boldsymbol{x}^{\prime}$ in Equation 5 may have both domain and semantic shifts. With the new PatchMix augmented embedding layer $\bar{\varphi}$ and the two projection heads of $\tilde{\mathcal{H}}$, we can obtain $[\bar{\boldsymbol{z}_{d}},\bar{\boldsymbol{z}_{s}}]=\tilde{\mathcal{H}}(\mathcal{F}(\bar{\varphi}(\boldsymbol{x})))$. We separately consider the learning of domain and semantic features. For semantic features, we introduce a factor $\alpha$ which takes the portion semantic of the sample $\boldsymbol{x}$ into account, after mixing up with $\boldsymbol{x}^{\prime}$. In specific, the contrastive loss in Equation 1 is now modified as:
$\mathcal{L}^{rep}_{s}(\boldsymbol{x})=-\frac{1}{|\mathcal{P}(\boldsymbol{x})|}\sum_{\boldsymbol{\bar{z}}^{+}_{s}\in\mathcal{P}(\boldsymbol{x})}\alpha\log\sigma(\boldsymbol{\bar{z}}_{s}\cdot\boldsymbol{\bar{z}}^{+}_{s};\tau),$ (6)
@@
where $\alpha=\frac{\boldsymbol{\beta}\cdot{\boldsymbol{s}}}{\boldsymbol{\beta}\cdot\boldsymbol{s}+(1-\boldsymbol{\beta})\cdot\boldsymbol{s}^{\prime}}$. $\boldsymbol{\beta}$ denotes the vector consisting of all $\beta_{j}$ as in Equation 5. $\boldsymbol{s}$ and $\boldsymbol{s}^{\prime}$ are two vectors storing the attention scores for all the patches for $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ respectively. The attention scores, computed following [74, 17], account for the semantic weight of each patch. To train the semantic classification head, we adopt the loss as in Equation 2. Differently, if $\boldsymbol{x}$ is a labelled sample, inspired by [75], we replace $\boldsymbol{q}$ with $\boldsymbol{\bar{q}}=\alpha\cdot\boldsymbol{q}+\frac{1-\alpha}{|\mathcal{C}|}\cdot\boldsymbol{1}$, where $\boldsymbol{q}$ is the one-hot vector derived from the label $y$ of $\boldsymbol{x}$. If $\boldsymbol{x}$ is unlabelled, similar to Equation 2, $\boldsymbol{q}$ is a pseudo-label from a sharpened prediction of from another mixed-up view. Aside from the label $\boldsymbol{q}$, we also need to learn another set of semantic prototypes by replacing $\boldsymbol{W}$ with $\boldsymbol{W}^{s}$. Let the modified classification loss be $\mathcal{L}^{cls}_{s}$ and $\mathcal{L}^{r,c}_{s}$ be the summation of $\mathcal{L}^{rep}_{s}$ and $\mathcal{L}^{cls}_{s}$. The loss for PatchMix-based semantic representation and classification learning is:
$\mathcal{L}_{s}=\lambda\sum_{\boldsymbol{x}\in\mathcal{B}}\mathcal{L}^{r,c}_{s}(\boldsymbol{x})+(1-\lambda)\sum_{\boldsymbol{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}_{s}(\boldsymbol{x}).$ (7)
@@
Next, for domain-specific features, we employ the same loss as Equation 6, except that we now train on $\bar{\boldsymbol{z}}_{d}$, for representation learning. We denote this loss as $\mathcal{L}^{rep}_{d}$. For training the classification head, we again adopt the loss as in Equation 2 but modify the label $\boldsymbol{q}$ and learn a set of domain prototypes $\boldsymbol{W}^{d}$. Therefore, if $\boldsymbol{x}$ is a labelled sample, $\boldsymbol{\bar{q}}=\alpha\cdot\boldsymbol{q}+\frac{1-\alpha}{\left|\Omega\right|}\cdot\boldsymbol{1}$, where $\boldsymbol{q}$ is the domain label. Note that we only assume that the labelled samples are from the same domain and do not assume that domain labels are available for any unlabelled samples, which is more realistic and challenging. Therefore, the only known domain label is typically 1. To obtain pseudo-labels for the unlabelled samples, we run the semi-supervised k-means as in [8] on the current mini-batch. We denote this modified classification loss as $\mathcal{L}^{cls}_{d}$ and the summation of $\mathcal{L}^{rep}_{d}$ and $\mathcal{L}^{cls}_{d}$ as $\mathcal{L}^{r,c}_{d}$. Therefore, the loss for representation and classification learning can be written as
$\mathcal{L}_{d}=\lambda\sum_{\boldsymbol{x}\in\mathcal{B}}\mathcal{L}^{r,c}_{d}(\boldsymbol{x})+(1-\lambda)\sum_{\boldsymbol{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}_{d}(\boldsymbol{x}).$ (8)
@@ 3.2.3 Curriculum sampling
As curriculum sampling [26] can effectively enhance the generalization capability of models by gradually increasing the difficulty of the training data, which is also a natural fit to the GCD with domain shift problem. Here, we also introduce a curriculum sampling scheme to further enhance the learning of our HiLo framework. We expect the training to start by focusing on samples from the same domain to learn semantic features and leverage more samples containing the additional challenge of domain shifts in the later training stages. To this end, we devise a difficulty measure $p_{cs}(\boldsymbol{x}|t)$ for each sample $\boldsymbol{x}$ at training time step $t$ (i.e., epoch), by considering the portion of samples belonging to each domain. As the unlabelled samples are from multiple domains and we do not have access to the domain label, we run the semi-supervised k-means on all the domain features $\boldsymbol{z}_{d}$ extracted using the DINO pretrained backbone. Let the resulting clusters along the domain axis be $\hat{\mathcal{D}}^{a}$ and $\hat{\mathcal{D}}^{b}$, which corresponds to domains $\Omega^{a}$ and $\Omega^{b}$ respectively and $\mathcal{D}^{u}=\hat{\mathcal{D}}^{a}\cup\hat{\mathcal{D}}^{b}$. With the above, we then define the sampling probability weight $p_{cs}(\boldsymbol{x}|t)$ for each sample as follows:
@@
$p_{cs}(\boldsymbol{x}|t)=\begin{dcases}\begin{aligned} &\phantom{-}\phantom{-}\phantom{-}\phantom{-}1,\qquad\qquad\qquad\quad\boldsymbol{x}\in\mathcal{D}^{l}\\ &\phantom{-}\phantom{-}\phantom{-}\frac{|\mathcal{D}^{l}|}{|\hat{\mathcal{D}}^{a}|},\qquad\qquad\qquad\ \boldsymbol{x}\in\hat{\mathcal{D}}^{a},\\ &r_{0}+(r^{\prime}-r_{0})\mathds{1}(t>t^{\prime}),\quad\quad\!\!\boldsymbol{x}\in\hat{\mathcal{D}}^{b}\\ \end{aligned}\end{dcases}$ (10)
@@
where $\mathds{1}(\cdot)$ is an indicator function, $t^{\prime}$ is a constant epoch number since which we would like to increase the portion of samples from unknown domains, $r_{0}$ and $r^{\prime}$ are constant probabilities for samples from unknown domains to be sampled in the earlier stages (i.e., $<t^{\prime}$) and latter stages (i.e., $>t^{\prime}$), $t$ indicates the current training time step. In our formulation, (1) if $\boldsymbol{x}$ is a labelled sample, its $p_{cs}(\boldsymbol{x}|t)$ is set to 1, without any discount; (2) if $\boldsymbol{x}$ is an unlabelled sample and is in $\hat{\mathcal{D}}^{a}$ (i.e., predicted as from the seen domain), $p_{cs}(\boldsymbol{x}|t)$ is set to $\frac{|\mathcal{D}^{l}|}{|\hat{\mathcal{D}}^{a}|}$ (i.e., proportional to the labelled and unlabelled samples from the same domain, as per the sampling strategy used in the conventional GCD without domain shifts [8]); and (3) if $\boldsymbol{x}$ is an unlabelled sample and is in $\hat{\mathcal{D}}^{a}$ (i.e., predicted as from the unseen domain), its $p_{cs}(\boldsymbol{x}|t)$ will increase along with the training after epoch $t^{\prime}$.
@@
In Appendix D, we provide an approximated theoretical analysis for our method. Theorem 1 suggests (1) that learning on the original domain data first can effectively lower the error bound of category discovery on $\mathcal{D}^{u}$ and (2) the domain head that can reliably discriminate original and new domain samples can further reduce this error bound. Theorem 2 suggests that minimizing the mutual information between domain and semantic features can further lower the error bound of category discovery on $\mathcal{D}^{u}$. These theorems further validate the effectiveness of our method from a theoretical perspective.

### v2

@@ 3.2.1 Learning domain-semantic disentangled features for GCD
As covariate shift observed by new domains $\Omega$ in $\mathcal{D}^{u}$ degrades performance, we aim to learn two distinct feature sets encoding domain and semantic aspects by minimizing their mutual information. For each image $\bm{x}$, we thus consider that its feature can be partitioned into two parts, depicting domain-specific (e.g., real, sketch) and semantic information (e.g., cat, dog), respectively. However, it is intractable to estimate the mutual information between random variables of semantic and domain in finite high-dimensional space without parametric assumptions Zhao et al. (2018); Song & Ermon (2019). Instead of calculating the exact value, assumptions based on convex conjugate Nguyen et al. (2010) and GAN Nowozin et al. (2016) are utilized for estimation. Belghazi et al. (2018); Hjelm et al. (2018) further demonstrate that this estimation can be achieved without such assumptions. We thus adopt the approach from Hjelm et al. (2018) based on Jensen-Shannon divergence to estimate the mutual information. For each image, instead of considering a single feature vector as $\bm{z}=\mathcal{H}(\mathcal{F}(\varphi(\bm{x})))$, here we consider two feature vectors, $\bm{z}_{d}$ and $\bm{z}_{s}$, for domain and semantic information respectively. Inspired by the fact that deeper layers of the model give higher-level features and the shallower layers of the model give lower-level features Sze et al. (2017); Zhou et al. (2021), we use the feature from the very first layer of the ViT as $\bm{z}_{d}$ and that from the very last layer as $\bm{z}_{s}$. Specifically, we obtain $[\bm{z}_{d},\bm{z}_{s}]=\tilde{\mathcal{H}}(\mathcal{F}(\varphi(\bm{x})))$, where $\tilde{\mathcal{H}}$ consists of two projection heads, one on the first layer feature of $\mathcal{F}$ and the other on the last layer feature of $\mathcal{F}$ (see Figure 2). Therefore, the mutual information between domain and semantic features can be approximated by a Jensen-Shannon estimator:
$\mathcal{L}_{m}={I}_{\Phi}(\bm{z}_{d},\bm{z}_{s})=\mathbb{E}_{p(\bm{z}_{d},\bm{z}_{s})}\left[-\log\left(1+e^{-\Phi(\bm{z}_{d},\bm{z}_{s})}\right)\right]-\mathbb{E}_{p(\bm{z}_{d})p(\bm{z}_{s})}\left[\log\left(1+e^{\Phi(\bm{z}_{d},\bm{z}_{s})}\right)\right],$ (4)
where $\Phi$ is an MLP and an output dimension of 1. $\Phi$ takes the concatenation of $\bm{z}_{s}$ and $\bm{z}_{d}$ as input and predict a single scalar value. We aim to minimize the expected log-ratio of the joint distribution concerning the product of marginals. Note that here $\bm{z}_{s}$ and $\bm{z}_{d}$ may come from two different images. In practice, we tile the domain and semantic features of all the images in the mini-batch, and concatenate them, before applying $\Phi$ on all the concatenated features. We then extract the diagonal entries (which are from the marginals) as the first term and the other entries (which are from the joint distribution) as the second term in Equation 4.
@@ 3.2.2 PatchMix contrastive learning
Mixup Zhang et al. (2018b) is a powerful data augmentation technique that involves blending pairs of samples and their corresponding labels to create new synthetic training examples. It has been shown to be very effective in semi-supervised learning Hataya & Nakayama (2019), long-tailed recognition Xu et al. (2021), etc. In the presence of domain shifts, Mixup has also been shown to be effective in unsupervised domain adaptation Na et al. (2021) and domain generalization Zhang et al. (2018a); Yun et al. (2019); Zhou et al. (2021). Recently, PMTrans Zhu et al. (2023) introduced PatchMix, which is a variant of Mixup augmentation by mixing up the embeddings of images in the Transformer-based architecture for domain adaptation. Particularly, for an input image $\bm{x}$ with label $y$, PatchMix augments its $j$-th embedding patch by
$\bar{\varphi}(\bm{x})_{j}={\beta}_{j}\odot\varphi(\bm{x})_{j}+(1-{\beta}_{j})\odot\varphi(\bm{x}^{\prime})_{j},$ (5)
@@
where $\bm{x}^{\prime}$ is an unlabelled image with or without domain shift, ${\beta}_{j}\in[0,1]$ is the random mixing proportion for the $j$-th patch, sampled from Beta distribution, and $\odot$ denotes the multiplication operation. A one-hot vector derived from $y$ is then smoothed based on $\beta_{j}$ to supervise the cross-entropy loss to train the classification model. However, this works under the assumption that the out-of-domain samples share the same class space with the in-domain samples, restricting its application to the more practical scenarios where the out-of-domain samples may come from new classes as we consider in the problem of GCD with domain shift. Hence, we devise a PatchMix-based contrastive learning method to address the challenge of GCD in the presence of domain shift. Our approach properly leverages all available samples, including both labelled and unlabelled data, from both in-domain and out-of-domain sources, encompassing both old and new classes. By incorporating these diverse samples, our technique aims to improve the model’s ability to handle domain shifts and effectively generalize across different classes.
@@
When incorporating the PatchMix into our problem setting, the unlabelled sample $\bm{x}^{\prime}$ in Equation 5 may have both domain and semantic shifts. With the new PatchMix augmented embedding layer $\bar{\varphi}$ and the two projection heads of $\tilde{\mathcal{H}}$, we can obtain $[\bar{\bm{z}_{d}},\bar{\bm{z}_{s}}]=\tilde{\mathcal{H}}(\mathcal{F}(\bar{\varphi}(\bm{x})))$. We separately consider the learning of domain and semantic features. For semantic features, we introduce a factor $\alpha$ which takes the portion semantic of the sample $\bm{x}$ into account, after mixing up with $\bm{x}^{\prime}$. In specific, the contrastive loss in Equation 1 is now modified as:
$\mathcal{L}^{rep}_{s}(\bm{x})=-\frac{1}{|\mathcal{P}(\bm{x})|}\sum_{\bm{\bar{z}}^{+}_{s}\in\mathcal{P}(\bm{x})}\alpha\log\sigma(\bm{\bar{z}}_{s}\cdot\bm{\bar{z}}^{+}_{s};\tau),$ (6)
@@
where $\alpha=\frac{\bm{\beta}\cdot{\bm{s}}}{\bm{\beta}\cdot\bm{s}+(1-\bm{\beta})\cdot\bm{s}^{\prime}}$. $\bm{\beta}$ denotes the vector consisting of all $\beta_{j}$ as in Equation 5. $\bm{s}$ and $\bm{s}^{\prime}$ are two vectors storing the attention scores for all the patches for $\bm{x}$ and $\bm{x}^{\prime}$ respectively. The attention scores, computed following Chen et al. (2022a); Zhu et al. (2023), account for the semantic weight of each patch. To train the semantic classification head, we adopt the loss as in Equation 2. Differently, if $\bm{x}$ is a labelled sample, inspired by Szegedy et al. (2016), we replace $\bm{q}$ with $\bm{\bar{q}}=\alpha\cdot\bm{q}+\frac{1-\alpha}{|\mathcal{C}|}\cdot\bm{1}$, where $\bm{q}$ is the one-hot vector derived from the label $y$ of $\bm{x}$. If $\bm{x}$ is unlabelled, similar to Equation 2, $\bm{q}$ is a pseudo-label from a sharpened prediction of from another mixed-up view. Aside from the label $\bm{q}$, we also need to learn another set of semantic prototypes by replacing $\bm{W}$ with $\bm{W}^{s}$. Let the modified classification loss be $\mathcal{L}^{cls}_{s}$ and $\mathcal{L}^{r,c}_{s}$ be the summation of $\mathcal{L}^{rep}_{s}$ and $\mathcal{L}^{cls}_{s}$. The loss for PatchMix-based semantic representation and classification learning is:
$\mathcal{L}_{s}=\lambda\sum_{\bm{x}\in\mathcal{B}}\mathcal{L}^{r,c}_{s}(\bm{x})+(1-\lambda)\sum_{\bm{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}_{s}(\bm{x}).$ (7)
@@
Next, for domain-specific features, we employ the same loss as Equation 6, except that we now train on $\bar{\bm{z}}_{d}$, for representation learning. We denote this loss as $\mathcal{L}^{rep}_{d}$. For training the classification head, we again adopt the loss as in Equation 2 but modify the label $\bm{q}$ and learn a set of domain prototypes $\bm{W}^{d}$. Therefore, if $\bm{x}$ is a labelled sample, $\bm{\bar{q}}=\alpha\cdot\bm{q}+\frac{1-\alpha}{\left|\Omega\right|}\cdot\bm{1}$, where $\bm{q}$ is the domain label. Note that we only assume that the labelled samples are from the same domain and do not assume that domain labels are available for any unlabelled samples, which is more realistic and challenging. Therefore, the only known domain label is typically 1. To obtain pseudo-labels for the unlabelled samples, we run the semi-supervised k-means as in Vaze et al. (2022) on the current mini-batch. We denote this modified classification loss as $\mathcal{L}^{cls}_{d}$ and the summation of $\mathcal{L}^{rep}_{d}$ and $\mathcal{L}^{cls}_{d}$ as $\mathcal{L}^{r,c}_{d}$. Therefore, the loss for representation and classification learning can be written as
$\mathcal{L}_{d}=\lambda\sum_{\bm{x}\in\mathcal{B}}\mathcal{L}^{r,c}_{d}(\bm{x})+(1-\lambda)\sum_{\bm{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}_{d}(\bm{x}).$ (8)
@@ 3.2.3 Curriculum sampling
As curriculum sampling Bengio et al. (2015) can effectively enhance the generalization capability of models by gradually increasing the difficulty of the training data, which is also a natural fit to the GCD with domain shift problem. Here, we also introduce a curriculum sampling scheme to further enhance the learning of our HiLo framework. We expect the training to start by focusing on samples from the same domain to learn semantic features and leverage more samples containing the additional challenge of domain shifts in the later training stages. To this end, we devise a difficulty measure $p_{cs}(\bm{x}|t)$ for each sample $\bm{x}$ at training time step $t$ (i.e., epoch), by considering the portion of samples belonging to each domain. As the unlabelled samples are from multiple domains and we do not have access to the domain label, we run the semi-supervised k-means on all the domain features extracted using the DINO pretrained backbone. Let the resulting clusters along the domain axis be $\hat{\mathcal{D}}^{a}$ and $\hat{\mathcal{D}}^{b}$, which corresponds to domains $\Omega^{a}$ and $\Omega^{b}$ respectively and $\mathcal{D}^{u}=\hat{\mathcal{D}}^{a}\cup\hat{\mathcal{D}}^{b}$. With the above, we then define the sampling probability weight $p_{cs}(\bm{x}|t)$ for each sample as follows:
@@
$p_{cs}(\bm{x}|t)=\begin{dcases}\begin{aligned} &\phantom{-}\phantom{-}\phantom{-}\phantom{-}1,\qquad\qquad\qquad\quad\bm{x}\in\mathcal{D}^{l}\\ &\phantom{-}\phantom{-}\phantom{-}\frac{|\mathcal{D}^{l}|}{|\hat{\mathcal{D}}^{a}|},\qquad\qquad\qquad\ \bm{x}\in\hat{\mathcal{D}}^{a},\\ &r_{0}+(r^{\prime}-r_{0})\mathds{1}(t>t^{\prime}),\quad\quad\!\!\bm{x}\in\hat{\mathcal{D}}^{b}\\ \end{aligned}\end{dcases}$ (10)
@@
where $\mathds{1}(\cdot)$ is an indicator function, $t^{\prime}$ is a constant epoch number since which we would like to increase the portion of samples from unknown domains, $r_{0}$ and $r^{\prime}$ are constant probabilities for samples from unknown domains to be sampled in the earlier stages (i.e., $<t^{\prime}$) and latter stages (i.e., $>t^{\prime}$), $t$ indicates the current training time step. In our formulation, (1) if $\bm{x}$ is a labelled sample, its $p_{cs}(\bm{x}|t)$ is set to 1, without any discount; (2) if $\bm{x}$ is an unlabelled sample and is in $\hat{\mathcal{D}}^{a}$ (i.e., predicted as from the seen domain), $p_{cs}(\bm{x}|t)$ is set to $\frac{|\mathcal{D}^{l}|}{|\hat{\mathcal{D}}^{a}|}$ (i.e., proportional to the labelled and unlabelled samples from the same domain, as per the sampling strategy used in the conventional GCD without domain shifts Vaze et al. (2022)); and (3) if $\bm{x}$ is an unlabelled sample and is in $\hat{\mathcal{D}}^{a}$ (i.e., predicted as from the unseen domain), its $p_{cs}(\bm{x}|t)$ will increase along with the training after epoch $t^{\prime}$. We investigate choices of $r_{0}$, $r^{\prime}$ and $t^{\prime}$ in Appendix M.
@@
In Appendix D, we provide an approximated theoretical analysis for our method with two theorems. Theorem 1 suggests (1) that learning on the original domain data first can effectively lower the error bound of category discovery on $\mathcal{D}^{u}$ and (2) the domain head that can reliably discriminate original and new domain samples can further reduce this error bound. Theorem 2 suggests that minimizing the mutual information between domain and semantic features can further lower the error bound of category discovery on $\mathcal{D}^{u}$. These theorems further validate the effectiveness of our method from a theoretical perspective.